Matt is seven years older than his sister,Vanessa.The sum of their ages is thirty-nine. Which system of equations describes Matt’s (M) and Vanessa’s (V) ages?
Answer by jorel555(1290)

M+V=39
M=V+7
Solving:
V+V+7=39
2V=32
V=16
Vanessa is 16; Matt is 23!!!!!!!!!!!!!